# RT-GuIDE: Real-Time Gaussian splatting for Information-Driven Exploration

### v1

@@ Abstract
We propose a framework for active mapping and exploration that leverages Gaussian splatting for constructing information-rich maps. Further, we develop a parallelized motion planning algorithm that can exploit the Gaussian map for real-time navigation. The Gaussian map constructed onboard the robot is optimized for both photometric and geometric quality while enabling real-time situational awareness for autonomy. We show through simulation experiments that our method is competitive with approaches that use alternate information gain metrics, while being orders of magnitude faster to compute. In real-world experiments, our algorithm achieves better map quality (10% higher Peak Signal-to-Noise Ratio (PSNR) and 30% higher geometric reconstruction accuracy) than Gaussian maps constructed by traditional exploration baselines. Experiment videos and more details can be found on our project page: https://tyuezhan.github.io/RT_GuIDE/
@@ I Introduction
Fig. 1 shows the elements of our approach. We use the Gaussian splatting approach proposed in [14] to generate Gaussian maps. We propose a novel information gain metric to compute the utility of regions in the environment and use a hierarchical planning framework to plan high-level navigation targets that yield maximal information in the environment and low-level paths that are dynamically feasible, collision-free and maximize the local information of the path. Our system runs real-time onboard a fully autonomous unmanned ground vehicle (UGV) to explore an unknown environment while generating a high-fidelity visual representation of the environment.
Figure 1: Key elements of our proposed approach. [A] Robot building a Gaussian map onboard in real-time and using it to avoid obstacles in the environment. Synthesized color and depth images from the Gaussian map are presented next to the corresponding observations from the RGBD sensor. [B] Robot navigating to unobserved areas (right) with high information gain while maximizing information along the trajectory. Gaussian frontiers colored cyan (light to dark) in increasing estimated information gain.
@@
    A unified framework for online mapping and planning with Gaussians.
@@
    An approximate information gain metric that is easy to compute and capable of running in real-time onboard an unmanned ground vehicle.
@@
    A method to compute frontiers in a map built using Gaussian splatting.
  4. 4.
    Extensive experiments in both indoor and outdoor environments to evaluate our framework.
@@ II Related Work
Map Representation. To effectively construct a map of the environment, numerous map representations have been proposed in the robotics community. The most intuitive but effective volumetric representation has been widely used. Voxel-based representation could maintain information such as occupancy [15] or signed distance [16, 17]. With the recent application of semantic segmentation, semantic maps that contain actionable information have been proposed [18, 19, 20, 21, 22]. With the recent advances in learned map representations in the computer vision community, Neural Radiance Fields (NeRF) [4] and 3D Gaussian Splatting (3DGS) [5] have become popular representations for robotic motion planning. In this work, we study the problem of active mapping with the learned map representations.
@@
Active Simultaneous Localization and Mapping (SLAM). The problem of exploration and active SLAM has been widely studied in the past decade. The classical exploration framework uses a model-based approach to actively navigate towards frontiers [11, 12] or waypoints that have the highest information gain [6, 7, 8, 9, 10, 23, 24]. Some recent work combines the idea of frontier exploration and the information-driven approach to further improve efficiency [25, 26, 27, 28]. However, most of the existing work developed their approaches based on classical map representations. For example, frontiers are typically defined in voxel maps, information metrics are typically associated with occupancy maps or Signed Distance Field (SDF) maps.
@@
In this work, we instead consider the active mapping problem with a learned map representation. Bayesian neural networks [29, 30] and deep ensembles [24, 31] are common approaches for estimating uncertainty in learned representations. [32, 33] use the idea of computing the uncertainty based on consecutive measurements from a moving robot in contrast to train multiple models in the traditional ensemble frameworks. Radiance field representations provide additional possibilities for estimating uncertainty through the volumetric rendering process [34]. [35] uses the difference between rendered depth and observations, and the alignment of $\alpha$-blended and median depth as measures of uncertainty.  [36] leverages Fisher information to compute pixel-wise uncertainty on rendered images. Where prior work in implicit and radiance field representations necessitates use indirect methods like ensembles and rendering for estimating uncertainty in the representation, a Gaussian map representation encodes physical parameters of the scene, which motivates estimating information gain from the Gaussians directly.
@@
Navigation in Radiance Fields. Prior work has also considered planning directly in radiance fields. [37] plans trajectories in a Gaussian map and uses observability coverage and reconstruction loss stored in a voxel grid as an approximation of information gain for exploration. Sim-to-real approaches leverage the rendering quality of the learnt radiance fields to train downstream tasks such as visual localization, imitation learning [38] by effectively utilizing the representation as a simulator. [39] use a liquid neural network to train imitation policies for a quadrotor robot in a Gaussian splat. [40] utilize the geometric fidelity of the representation to first map the environment and then perform trajectory optimization to plan paths in these environments. [41] use a pre-generated Gaussian map to compute safe polytopes to generate paths and also present a method to synthesize novel views with a coarse localization of the robot and refine the estimate by solving a Perspective-n-Point (PnP) problem. The probabilistic representations of free space has also been utilized for motion planning [42] where authors use uncertainties in the learned representation to provide probabilistic guarantees on collision-free paths.
III Problem Specification
@@
Figure 2: The proposed active mapping framework. The proposed framework contains two major components, the planning module and the mapping module. As can be seen in the figure, the Mapping module ([A]) takes in RGB, depth and pose measurements, and updates the map representation $m$ at every step and computes the utility of geometrically clustered frontiers. The information is then passed to the planning module which comprises the topological graph and motion primitive library ([B]). The topological graph adds sampled viewpoints as nodes and passes along a planned path to the trajectory planner. The trajectory planner in turn attempts to plan a path to goal that maximizes information gathering (queried from the mapper). The planned trajectory is executed by the robot to get a new set of observations.
We have an unknown map of the environment $m^{*}$ that belongs to a space of maps $\mathcal{M}$. Our goal is to solve the following problem
@@
Information-Driven Exploration.
$\min_{x_{1:T}}$ $\ \mathbb{E}_{x}[\rho(h(x,\hat{m}),h(x,m^{*}))]$ (1)
$s.t$ $\ \hat{m}=\Phi(y_{1:T},x_{1:T})$
$\ y_{t}=h(x_{t},m^{*})\ \forall 1\leq t\leq T.$
In the equation above, the expectation is taken over a random evaluation point $x$. The states belong to the state space $\mathcal{X}$, whereas observations belong to the observation space $\mathcal{Y}$. The observation function $h:\mathcal{X}\times\mathcal{M}\rightarrow\mathcal{Y}$ takes as input the robot state together with the map of the environment, and outputs the rendered image. $T$ represents the exploration budget. $m^{*}$ is the true map of the environment. $\hat{m}$ is the map of the environment that is estimated from the sequence of measurements $(y_{t})_{1\leq t\leq T}$ taken from vantage points (states) $(x_{t})_{1\leq t\leq T}.$ The discrepancy function $\rho$ captures the difference between the image rendered based on the estimated map, and the actual map. The sequence of vantage points is constrained to lie in free space $\mathcal{X}_{free}$.
Since Prob (1) is parameterized by $m^{*}$, which is a priori unknown, the problem is ill-posed. We are instead interested in synthesizing a (near) optimal policy $(\pi_{t})_{t=0}^{T}$, consisting of a sequence of functions that map the history of observations to the optimal action to take at the current step. In particular, for any $i\leq T$, the function
$\pi_{t}:\mathcal{Y}_{1}\times\cdots\times\mathcal{Y}_{t}$ $\rightarrow\mathcal{X}$ (2)
$\pi_{t}(y_{1},y_{2},...,y_{t})$ $=x_{t}$
ought to specify the next vantage point given the the history of observations accrued thus far. Finding the optimal sequence of policies is a challenging, high-dimensional, optimization problem. We therefore develop an approximate method for solving it. The approximation lies in a particular, but suitable, choice of representation for $m^{*}$, together with a way of using it for effective exploration.
IV Method
Our proposed framework comprises the mapping module and the planning module, as illustrated in Fig. 2. The mapping module (Sec. IV-A) accumulates measurements and poses to generate a reconstruction of the environment and computes the utility of Gaussian frontiers. The information is then passed to the planning module for planning guidance paths (Sec. IV-B1) and trajectories (Sec. IV-B2).
IV-A Mapping & Uncertainty Estimation
IV-A1 Mapping
We adopt the 3D Gaussian Splatting (3DGS) approach to represent the environment as a map of 3D Gaussians. 3DGS enables real-time radiance field rendering by projecting and blending the 3D Gaussians to render images. The parameters of the 3D Gaussians are optimized through mapping iterations to improve the representation of the scene. We build the 3DGS mapping module upon the [14] framework. The poses of the robot are obtained using [43].
IV-A2 Gaussian uncertainty for viewpoint selection
We estimate uncertainty in the map representation directly on the 3D Gaussians. We motivate our metric of information gain leveraging the theory of Kalman filtering. In particular, the updates of the Kalman filter [44] are given by
$m_{k|k}$ $=m_{k|k-1}+K(y_{k}-Hx_{k|k-1})$ (3)
$K$ $=P_{k|k-1}H^{T}S^{-1}$
$S$ $=HP_{k|k-1}H^{T}+R$
$P_{k|k}$ $=P_{k|k-1}-KHP_{k|k-1}.$
where $K$ is the Kalman Gain, $P$ is the covariance of the map, $H$ is the linearized measurement model, $S$ is the innovation covariance and $R$ is the covariance of the measurement noise. The following identity holds:
$Cov_{k|k-1}(m_{k|k}-m_{k|k-1})=P_{k|k-1}-P_{k|k}$ (4)
The pertinent aspect of the relation above is that the prior covariance of the change in the map parameters upon observing a new measurement is equal to the information gain of the latter. Although this argument can be rigorously shown in the context of linear measurement models, we nevertheless use it as a measure of future information gain in the context of measurements with a Gaussian representation of the environment.
Gaussian uncertainty. Following the above motivation, we compute the change in Gaussian parameters across each mapping optimization and use the magnitude of the parameter updates $\|\hat{m}_{t}-\hat{m}_{t-1}\|$ as a measure of uncertainty. Gaussians with large changes in parameters represent parts of the scene with high uncertainty and conversely Gaussians with stable parameters have low uncertainty. We use this to approximate information gain for viewpoint selection.
Given a map $m$ represented by a set of Gaussians $G$, we classify the Gaussians in the scene $G=\{G_{H},G_{L},G_{O}\}$ into three disjoint subsets, $G_{H}$ which consists of Gaussians with high uncertainty, $G_{L}$ which consists of Gaussians with low uncertainty and $G_{O}$ for the rest. For a Gaussian $g_{i}\in G$ and pose $x_{t}$ at timestep $t$, we define the binary visibility function $v(x_{t},g_{i})$ capturing whether $g_{i}$ is in the field of view of $x_{t}$.
To maximize information gain while exploring the environment, we aim to find the viewpoint with the highest utility $\xi(x_{t})$ that maximizes the number of high-uncertainty Gaussians and minimizes the number of low-uncertainty Gaussians in the field of view at each timestep. The first term in Equ. 5 encourages observations of parts of the scene with high uncertainty while the second term encourages exploration of unseen areas of the scene. The balance of exploitation and exploration is weighted by the term $\lambda_{\xi}$.
$\xi(x_{t})=\sum_{i\in L}v(x_{t},g_{i})-\lambda_{\xi}\sum_{i\in H}v(x_{t},g_{i})$ (5)
Gaussian frontiers. In contrast to traditional mapping representations, the Gaussian map does not encode occluded and free space. Instead of frontiers of unobserved regions, we use the Gaussian uncertainty estimates to identify regions of the map that should be visited next, since Gaussians that have few observations or lie at the edges of observed regions naturally have higher uncertainty. For a region $o\subset m$, we consider the set of Gaussians $\{g\sim\mathcal{N}(\mu,\sigma^{2})\in G\ |\ \mu\in o\}$ with cardinality $N_{g}$ in the region and compute the mean uncertainty of that frontier
$\Omega=\frac{1}{N_{g}}\sum^{N_{g}}_{i}\|\mu^{t}_{i}-\mu^{t-1}_{i}\|.$ (6)
This provides a unified framework for estimating information gain to optimize map quality while implicitly identifying frontiers to guide the exploration of unknown regions.
IV-A3 Traversable region segmentation
We found Gaussians on the ground plane to be particularly noisy in indoor environments due to reflections off the floor. This necessitated the removal of Gaussians on the ground for planning. This also fulfilled an important role in reducing the number of Gaussians that have to be queried in the collision check.
IV-B Hierarchical Planning
We represent the planning problem as a hierarchical planning problem. The first level provides guidance to map particular regions of the environment and a path to that region from the known space in the environment. The second level finds a trajectory that is dynamically-feasible (i.e. obeys the robot’s physical constraints); collision-free and locally maximizes the information along the path.
IV-B1 High-level guidance
We formulate a high-level guidance path that allows us to (i) plan to transition to high uncertainty regions of space which might involve traveling through those with low uncertainty; and (ii) estimate the traversable space in the known space without explicitly having to plan long-range trajectories with computationally expensive collision checks. At the higher level of our planner, we construct a graph $\mathcal{G}(N,E)$ by incrementally adding nodes along the traveled path. The graph consists of two types of nodes $n_{o},n_{p}$ where $n_{o}$ are observed nodes (along our odometry) and $n_{p}$ are predicted nodes. At each planning iteration, we sample a fixed number of viewpoints around the identified regions. We can compute the shortest viewing distance from the Gaussian frontier $f\in\mathbb{R}^{3}$ to the optical center of the camera as
$d_{1}=\frac{||f^{T}\left[b_{1}\ b_{2}\right]||_{2}}{tan\left(\phi/2\right)}$
where $b$ is the camera axis and $\phi$ is the field of view of the camera. Each sampled viewpoint is then assigned the utility computed by the Gaussian frontiers and connected to the closest point in the graph. We then use Dijkstra’s search algorithm to find the shortest path from the current robot location in the graph to all the predicted nodes in the graph. We then compute the cost-benefit of a path using $\Omega/e^{d_{2}}$ where $\Omega$ is the estimated information gain and $d_{2}$ is the distance to the node [45]. The maximal cost-benefit path is sent to the trajectory planner.
IV-B2 Trajectory Generation
The second level planning problem involves finding a collision-free trajectory from start to goal that meets the dynamic constraints of the robot while optimizing the information along the path. The problem is defined as follows:
Problem 2.
@@ IV-B2 Trajectory Generation
Trajectory Planning. Given an initial robot state $x_{0}\in\mathcal{X}_{free}$, and goal region $\mathcal{X}_{goal}$, find the control inputs $u(\cdot)$ defined on $[0,\tau]$ that solve:
$\min_{u(\cdot),\tau}\ J(u(\cdot),\tau)$ (7)
$s.t.\ \text{for all }t\in[0,\tau]$

### v2

@@ Abstract
We propose a framework for active mapping and exploration that leverages Gaussian splatting for constructing dense maps. Further, we develop a GPU-accelerated motion planning algorithm that can exploit the Gaussian map for real-time navigation. The Gaussian map constructed onboard the robot is optimized for both photometric and geometric quality while enabling real-time situational awareness for autonomy. We show through simulation experiments that our method yields comparable Peak Signal-to-Noise Ratio (PSNR) and similar reconstruction error to state-of-the-art approaches, while being orders of magnitude faster to compute. In real-world experiments, our algorithm achieves better map quality (at least 0.8dB higher PSNR and more than 16% higher geometric reconstruction accuracy) than maps constructed by a state-of-the-art method, enabling semantic segmentation using off-the-shelf open-set models. Experiment videos and more details can be found on our project page: https://tyuezhan.github.io/RT_GuIDE/
@@ I Introduction
Fig. 1 shows the elements of our approach. We use the Gaussian splatting approach proposed in [14] to generate maps. We propose an efficient information gain metric and use a hierarchical planning framework to plan high-level navigation targets that yield maximal information in the environment and low-level paths that are dynamically feasible, collision-free and maximize the local information of the path. Our system runs real-time onboard a fully autonomous unmanned ground vehicle (UGV) to explore an unknown environment while generating a high-fidelity visual representation of the environment.
Figure 1: Key elements of our proposed approach. [A] Robot building a Gaussian map onboard in real-time and using it to avoid obstacles in the environment. Synthesized color and depth images from the Gaussian map are presented next to the corresponding observations from the RGBD sensor. [B] Robot navigating to unobserved areas (right) with high information gain while maximizing information along the trajectory. Map regions colored light to dark in increasing estimated information gain.
@@
    A unified framework for online mapping and planning built only on a Gaussian splatting-based map, eliminating the need for additional volumetric representations.
@@
    An information gain heuristic that is easy to compute and capable of running in real-time onboard a robot.
@@
    A real-time exploration system built on Gaussian splatting, comprehensively validated in both simulation and real-world experiments across diverse indoor and outdoor environments. We will release the full autonomy framework as open-source.
@@ II Related Work
To effectively construct a map of the environment, numerous map representations have been proposed in the robotics community. The most intuitive but effective volumetric representation has been widely used. Voxel-based representation could maintain information such as occupancy [15] or signed distance [16, 17]. With the recent application of semantic segmentation, semantic maps that contain actionable information have been proposed [18, 19, 20, 21, 22]. With the recent advances in learned map representations in the computer vision community, Neural Radiance Fields (NeRF) [4] and 3D Gaussian Splatting (3DGS) [5] have become popular representations for robotic motion planning. In this work, we study the problem of active mapping with the learned map representations.
@@
The problem of exploration and active Simultaneous Localization and Mapping (SLAM) has been widely studied in the past decade. The classical exploration framework uses a model-based approach to actively navigate towards frontiers [11, 12] or waypoints that have the highest information gain [6, 7, 8, 9, 10, 23, 24]. Some recent work combines the idea of frontier exploration and the information-driven approach to further improve efficiency [25, 26, 27, 28]. However, most of the existing work developed their approaches based on classical map representations. For example, frontiers are typically defined in voxel maps, information metrics are typically associated with occupancy maps or Signed Distance Field (SDF) maps.
@@
In this work, we instead consider the active mapping problem with a learned map representation. Bayesian neural networks [29, 30] and deep ensembles [24, 31] are common approaches for estimating uncertainty in learned representations. [32, 33] use the idea of computing the uncertainty based on consecutive measurements from a moving robot in contrast to train multiple models in the traditional ensemble frameworks. Radiance field representations provide additional possibilities for estimating uncertainty through the volumetric rendering process [34]. [35] uses the difference between rendered depth and observations, and the alignment of $\alpha$-blended and median depth as measures of uncertainty.  [36] leverages Fisher information to compute pixel-wise uncertainty on rendered images. Where prior work in implicit and radiance field representations use indirect methods like ensembles and rendering for estimating uncertainty in the representation, a map represented by Gaussians encodes physical parameters of the scene, which motivates estimating information from the Gaussian parameters directly.
@@
Prior work has also considered planning directly in radiance fields. [37] plans trajectories in a Gaussian map and uses observability coverage and reconstruction loss stored in a voxel grid as an approximation of information gain for exploration. Sim-to-real approaches leverage the rendering quality of the learnt radiance fields to train downstream tasks such as visual localization, imitation learning [38] by effectively utilizing the representation as a simulator. [39] use a liquid neural network to train imitation policies for a quadrotor robot in a Gaussian splat. [40] utilize the geometric fidelity of the representation to first map the environment and then perform trajectory optimization to plan paths in these environments. [41] use a pre-generated Gaussian map to compute safe polytopes to generate paths and also present a method to synthesize novel views with a coarse localization of the robot and refine the estimate by solving a Perspective-n-Point (PnP) problem. The probabilistic representations of free space has also been utilized for motion planning [42] where authors use uncertainties in the learned representation to provide guarantees on collision-free paths.
III Problem Specification & Preliminaries
@@
Figure 2: The proposed active mapping framework. The proposed framework contains two major components, the planning module and the mapping module. As can be seen in the figure, the Mapping module ([A]) takes in RGB, depth and pose measurements, and updates the map representation at every step and computes the utility of cuboidal regions (Sec. IV-B1). The utility of each region is then passed to the planning module which comprises the topological graph and motion primitive library ([B]). Viewpoints are sampled around regions in the map and added to the topological graph. A high-level guidance path is generated from the topological graph to viewpoints with highest utility and passed to the trajectory planner. The trajectory planner in turn attempts to plan a path to goal that maximizes information gathering (queried from the mapper). The planned trajectory is executed by the robot to get a new set of observations.
Our goal is to construct an estimate $\hat{G}$ of the true map of the environment $G^{*}$, which is a priori unknown. The quality of the constructed map $\hat{G}$ is evaluated using a hidden test set $\mathcal{T}$ of tuples of poses and corresponding noise-free measurements, $(x_{test}^{*},y_{test}^{*})_{test\in\mathcal{T}}$. We aim to solve the following problem
$\mathop{\arg\min}\limits_{\hat{G},\ x_{1:T}}$ $\sum_{{test}\in\mathcal{T}}\mathcal{L}(y^{*}_{test},h(x^{*}_{test},\hat{G}))$ (1)
$s.t$ $\ \hat{G}=\Phi(x_{1:T},y_{1:T}),$
where $\mathcal{L}(\cdot)$ captures the difference between a synthesized and a true measurement, $\Phi(\cdot)$ is the function that constructs $\hat{G}$ from the set of measurements $y_{1:T}$ obtained at poses $x_{1:T}$, $h(\cdot)$ is the rendering function that synthesizes a measurement given a pose and a map, and $T$ is the time budget for exploration.
The problem above is ill-posed, as the set of test poses and measurements $\mathcal{T}$ is hidden. We approach this challenge by solving the following two problems in a receding horizon scheme. At time $k\leq T$, we first solve a mapping problem that involves minimizing the difference between the observed measurements $y_{1:k}$ and synthesized measurements $\hat{y}_{1:k}$:
$\Phi(x_{1:k},y_{1:k}):=\mathop{\arg\min}\limits_{\hat{G}}\sum_{s=1}^{k}\mathcal{L}(y_{s},h(x_{s},\hat{G})).$ (2)
Then, given the posterior estimate of the map $\hat{G}$, we find the subsequent viewpoint $x_{k+1}$ that maximizes the mutual information $\mathcal{I}(\cdot)$ between the map $\hat{G}$ and the measurement $y_{k+1}$ observed from the corresponding pose:
$\max_{x_{k+1}}\ \mathcal{I}(\hat{G};y_{k+1}|x_{k+1}).$ (3)
Gaussian Splatting. We use 3D Gaussian Splatting (3DGS) [5] to represent the environment as a collection of isotropic Gaussians $\hat{G}$. Each Gaussian is parameterized with 8 values representing RGB intensities ($c$), 3D position ($\mu$), radius ($r$), and opacity ($\alpha$). To generate the measurement at a given camera pose of the estimated map $\hat{G}$, 3D Gaussians are sorted in increasing order of depth relative to the camera pose and are then splatted onto the image plane. In particular, each splatted Gaussian will have a projected 2D mean ($\mu_{2D}$) and radius ($r_{2D}$) on the image plane,
$\mu_{2D}=\frac{KR_{k}^{\intercal}(\mu-T_{k})}{d},\ r_{2D}=\frac{fr}{d},\ d=e_{3}^{\intercal}R_{k}^{\intercal}(\mu-T_{k}),$
where $K$ is the intrinsic matrix, $R_{k}$ and $T_{k}$ are the rotational and translational components of the camera pose in the world frame at time $k$, $f$ is the focal length, and $e_{3}=[0,0,1]^{\intercal}$.
For a given pixel $p$ in the image, its color $\mathcal{C(\cdot)}$ and depth $\mathcal{D(\cdot)}$ can be obtained as
$\mathcal{C}(p)=\sum_{i=1}^{n}c_{i}f_{i}(p)\prod_{j=1}^{i-1}(1-f_{j}(p))\$
and
$\mathcal{D}(p)=\sum_{i=1}^{n}d_{i}f_{i}(p)\prod_{j=1}^{i-1}(1-f_{j}(p)),$
where
$f_{i}(p)=\alpha_{i}\exp\left\{-\frac{||p-\mu_{2D,i}||^{2}}{2r_{2D,i}^{2}}\right\}.$
At every iteration, we render color image $\hat{y}_{c}$ and depth image $\hat{y}_{d}$. We set the function $\mathcal{L}$ (from eq. (2)) to
$\mathcal{L}$ $=\frac{1}{N_{p}}\sum_{p\in y}\left(|\hat{y}_{d}-y_{d}|+\lambda_{1}|\hat{y}_{c}-y_{c}|\right)$ (4)
$+\lambda_{2}(1-SSIM(\hat{y}_{c},y_{c}))$
which is a weighted combination of the L1 loss on depth and rendered pixel colors and the structured similarity index measure (SSIM) to update the parameters of the Gaussians, and $N_{p}$ is the number of pixels in the image.
IV Method
Our proposed framework comprises the mapping module and the planning module, as illustrated in Fig. 2. The mapping module (Sec. IV-A) accumulates measurements and poses to generate a reconstruction of the environment and computes the uncertainty of Gaussians in the map. This is then passed to the planning module for planning guidance paths (Sec. IV-B1) and trajectories (Sec. IV-B2).
IV-A Mapping & Uncertainty Estimation
We build the 3DGS mapping module upon the [14] framework. Similar to [14], we set all Gaussians to be isotropic. As in [14], we set $\lambda_{1}=0.4,\lambda_{2}=0.1$. At the beginning of each mapping iteration, we initialize new Gaussians at positions corresponding to the current color and depth measurements. The parameters of these Gaussians are refined through backpropagation. As in [5], we prune Gaussians with low opacity or large radius.
Uncertainty Estimation. The objective of finding the next-best-view can then be expressed as
$x^{*}_{k+1}=\mathop{\arg\max}\limits_{x_{k+1}}\mathcal{I}(y_{k+1};\hat{G}|x_{k+1}).$ (5)
The information gain $\mathcal{I}(\cdot)$ is defined as
$\mathcal{I}$ $(y_{{k+1}};\hat{G}|x_{k+1})$ (6)
$=\mathbb{H}(y_{k+1}|x_{k+1})-\mathbb{H}(y_{k+1}|\hat{G},x_{k+1})$
where $\mathbb{H}(\cdot)$ is the entropy. We make the following assumptions in this work to approximate the information gain efficiently: (a) $\hat{G}\sim\mathcal{N}(\mu_{\hat{G}},\Sigma_{\hat{G}})$; (b) $y_{k+1}|\hat{G}\sim\mathcal{N}(h(x_{k+1},\hat{G}),\Sigma_{y})$; (c) constant isotropic variance $\Sigma_{y}$. From (a) and (b), the information gain can be simplified as
$\mathcal{I}(y_{{k+1}};\hat{G}|x_{k+1})$ $=\frac{1}{2}\log\det\left(\frac{J_{k+1}\Sigma_{\hat{G}}J_{k+1}^{\intercal}+\Sigma_{y}}{\Sigma_{y}}\right)$ (7)
$=\frac{1}{2}\log\det(I+\Sigma_{y}^{-1}J_{k+1}\Sigma_{\hat{G}}J_{k+1}^{\intercal})$
where $J_{k+1}=\frac{\partial h(x_{k+1},\hat{G})}{\partial G}$. Using the first order approximation of $\log\det$ about the identity, we can define a proxy metric for scoring next best views as
$x^{*}_{k+1}=\mathop{\arg\max}\limits_{x_{k+1}}\mathrm{Tr}(\Sigma_{y}^{-1}J_{k+1}\Sigma_{\hat{G}}J_{k+1}^{\intercal}).$ (8)
As in [36], assumption (c) allows us to simplify the optimization by ignoring the effect of $\Sigma^{-1}_{y}$. Evaluating $J_{k+1}$ is computationally expensive because it typically requires rendering at new poses. To achieve real-time evaluation of multiple candidate viewpoints, we replace the Jacobian term $J_{k+1}$ with a binary matrix that corresponds to its sparsity pattern. Intuitively, this captures the set of Gaussians within the field of view at pose $x_{k+1}$. The key insight behind this approximation is that we estimate information gain using the current uncertainty, based on the assumption that local measurements lead to precise local map estimates. We show empirically that the exploration performance based on our metric (in Sec. V-A) is similar to the performance based on eq. (8). We propose a heuristic for estimating uncertainty $\Sigma_{\hat{G}}$ based on the magnitude of the change in means ($\mu$) of the Gaussians, associating larger displacements with higher uncertainty. The motivation behind this lies in the fact that $\Sigma_{\hat{G}}$ can be approximated with the empirical Fisher Information matrix [43], $\Sigma_{\hat{G}}\approx\nabla_{G}\mathcal{L}(x_{k},y_{k})\nabla_{G}\mathcal{L}(x_{k},y_{k})^{\intercal}$, which in turn can be directly linked to the square of the updates of Gaussian parameters tuned by the gradient descent algorithm. In practice, we observe that the square of the magnitude of the updates outlined above is sensitive to measurement noise (in Sec. V-A), so we use the magnitude of the change instead.
Empirically, we observe that the means of Gaussians located near boundaries between observed and unobserved space exhibit significant changes over successive updates, indicating high uncertainty, as visualized in Fig. 2. Consequently, directing exploration toward these high-uncertainty Gaussians is analogous to following frontiers in grid-based exploration, making this heuristic effective for autonomous exploration. Importantly, through careful bookkeeping, Gaussians that remain unchanged in the most recent update retain their previously computed displacement values, ensuring that our heuristic provides a consistent measure of each Gaussian’s movement over time. In addition to the boundaries between observed and unobserved space, areas within observed space that have insufficient measurements also exhibit relatively high uncertainty. This further directs the robot to explore these areas and gather more information, thereby improving the overall map quality.
IV-B Hierarchical Planning
In spite of the approximations introduced in the previous section to compute information gain obtained by navigating to a given area in space in a computationally efficient way, the task of finding such a region is still a challenging optimization problem. For this reason, we address the latter task through a hierarchical planning framework. The high-level planner provides guidance to map particular regions of the environment and a path to that region from the known space in the environment. The low-level (trajectory) planner finds a trajectory that is dynamically-feasible (i.e. obeys the robot’s physical constraints), collision-free and locally maximizes the information gain along the path.
IV-B1 High-level planner
In contrast to traditional mapping representations, the Gaussian map does not encode occluded and free space. Instead of computing (geometric) frontiers, we use the Gaussian uncertainty estimates to identify regions of the map that should be visited next. We evenly partition the a priori known enclosing space of the desired map into cuboidal regions. At time ${k}$, for a region $o$, denote the cardinality of the Gaussians in the region by $M_{k}$. We compute the mean uncertainty of that region $\Omega_{k}=1/{M_{k}}\left(\sum_{\mu\in o}\|\mu_{k}-\mu_{{k-1}}\|_{2}\right).$ We formulate a high-level guidance path that allows us to (i) navigate to regions of high uncertainty; and (ii) utilize the known traversable space to plan long-range trajectories without computationally expensive collision checks. At the high level of our planner, we construct a tree by incrementally adding nodes along the traveled path. The tree consists of two types of nodes: odometry nodes and viewpoint nodes. At each planning iteration, we sample a fixed number of viewpoints around the identified regions. We compute the shortest viewing distance from the Gaussian region centroid to the optical center of the camera, given the camera intrinsics. Each sampled viewpoint is then assigned the utility computed for that region and connected to the closest odometry node in the tree as a viewpoint node. We then use Dijkstra’s algorithm to find the shortest path from the current robot location in the tree to all the viewpoint nodes in the tree. Finally, we compute the cost-benefit of a path using $\Omega/e^{d}$ where $\Omega$ is the estimated information gain and $d$ is the distance to the node [44]. The maximal cost-benefit path is sent to the trajectory planner.
IV-B2 Trajectory planner
We partition the current map $\hat{G}$ into three disjoint subsets: $\hat{G}_{H}$, which consists of Gaussians with high uncertainty, $\hat{G}_{L}$, which consists of Gaussians with low uncertainty, and $\hat{G}_{O}$ for the rest. For a Gaussian $g\in\hat{G}$ and pose $x$, we define the binary visibility function $v(x,g)$ capturing whether $g$ is in the field of view of $x$. To avoid rendering each view to check for occlusions, we cull the Gaussians that are beyond the perception range of the sensor.
To maximize information gain while exploring the environment, we aim to obtain viewpoints with high utility $\xi$ that maximize the number of high-uncertainty Gaussians and minimize the number of low-uncertainty Gaussians in the field of view. Trajectories are evaluated based on the sum of the utilities of the viewpoints in each trajectory given by
$\xi(x,\hat{G})=\sum_{g\in\hat{G}_{H}}v(x,g)-\lambda_{\xi}\sum_{g\in\hat{G}_{L}}v(x,g).$ (9)
The first term in eq. (9) encourages additional observations of parts of the scene with high uncertainty. The second term penalizes observation of stable Gaussians in explored areas and encourages exploration of unseen parts of the environment. The two terms are weighted by the $\lambda_{\xi}$.
Given the path generated by the high-level planner, we select the furthest point $p_{goal}$ along the path that lies within a local planning range $R_{H}$, and set it as the center of the goal region $\mathcal{X}_{goal}$. We use the unicycle model as the robot dynamics $f(\cdot)$. The robot state $x=[p,\theta]\in\mathbb{R}^{2}\times S^{1}$ consists of its position ($p$) and heading ($\theta$). The control inputs $u=[v,\omega]\in\mathbb{R}^{2}$ consist of linear velocity ($v$) and angular velocity $(\omega)$. We solve the trajectory planning problem in 2 steps: (i) finding feasible, collision-free trajectory candidates, and (ii) selecting a trajectory that maximizes the information gain. The first problem is defined as follows:
@@ IV-B2 Trajectory planner
Trajectory Candidate Generation. Given an initial robot state $x_{0}\in\mathcal{X}_{free}$, and goal region $\mathcal{X}_{goal}$, find the control inputs $u(\cdot)$ defined on $[0,\tau]$ that solve:
[…]
$s.t.\ \forall t\in[0,\tau],\ x(0)=x_{0},\ x(\tau)\in\mathcal{X}_{goal},$